OpenStudy (anonymous):

Solve the following equation. (x - 3) /x = (x - 3) / (x - 6)

OpenStudy (saifoo.khan):

Cross multiply.

OpenStudy (anonymous):

Now divide both sides by (x - 3) (x - 6) = x O My God, no solution...

OpenStudy (anonymous):

@jiteshmeghwal9 please solve it further and tell what is the answer...

OpenStudy (anonymous):

@jiteshmeghwal9 How in the world did you get from the original equation to THAT? Looks like you multiplied the numerators to each other and the same for the denominators.

OpenStudy (anonymous):

That's not cross multiplication though!

OpenStudy (jiteshmeghwal9):

Oh! i m sorry.

OpenStudy (anonymous):

The original equation is: \[\frac{x-3}{x} = \frac{x-3}{x-6}\]

OpenStudy (jiteshmeghwal9):

(x-3)(x-6)=(x-3)(x) x^2-6x-3x+18=x^2-3x x^2-9x+18=x^2-3x x^2-x^2-9x+3x=-18 -6x=-18 x=-18/-6=3.

OpenStudy (jiteshmeghwal9):

Am i correct now.

OpenStudy (jiteshmeghwal9):

@Wired & @waterineyes

OpenStudy (anonymous):

Yes and no. While it solves your equation, put X into the original equation. You're dividing into zero.

OpenStudy (jiteshmeghwal9):

I m only giving the value of the ques.

OpenStudy (jiteshmeghwal9):

'x' in the ques

OpenStudy (anonymous):

In the first part of your equation you can just remove (x-3) from both sides, coming back to what was said earlier (now deleted) where 0=6. With your equation you get (0/3)=(0/-3) aka 3=-3. Another impossibility.

myininaya (myininaya):

\[\frac{x - 3}{x} =\frac{x -3 }{x - 6}\] So you guys crossed multiplied By the way x cannot be 0, x-6 cannot be 0 because if these are zeros then the expressions evaluated at those particular values for x then one of the sides will not exist while the other side has a value that does exist.

myininaya (myininaya):